# Tight Verification of Probabilistic Robustness in Bayesian Neural Networks

### 21 Jan 2024 v1

@@ Abstract
We introduce two algorithms for computing tight guarantees on the probabilistic robustness of Bayesian Neural Networks (BNNs). Computing robustness guarantees for BNNs is a significantly more challenging task than verifying the robustness of standard Neural Networks (NNs) because it requires searching the parameters’ space for safe weights. Moreover, tight and complete approaches for the verification of standard NNs, such as those based on Mixed-Integer Linear Programming (MILP), cannot be directly used for the verification of BNNs because of the polynomial terms resulting from the consecutive multiplication of variables encoding the weights. Our algorithms efficiently and effectively search the parameters’ space for safe weights by using iterative expansion and the network’s gradient and can be used with any verification algorithm of choice for BNNs. In addition to proving that our algorithms compute tighter bounds than the SoA, we also evaluate our algorithms against the SoA on standard benchmarks, such as MNIST and CIFAR10, showing that our algorithms compute bounds up to 40% tighter than the SoA.
@@ 1 Introduction
**footnotetext: Equal contribution.
@@ Acknowledgements
The authors are grateful to Peyman Hosseini for extensive comments on previous versions of this paper and help with the experiments. This work is partially supported by the DARPA Assured Autonomy programme (FA8750-18-C-0095), the “REXASI-PRO” H-EU project, call HORIZON-CL4-2021-HUMAN-01-01, Grant agreement ID: 101070028, the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence, and the UK Royal Academy of Engineering (CiET17/18-26).
@@ Appendix A Limitations
1    pW_0 = compute_interval_probs_weight(np.asarray(vW_0), marg=w_margin, mean=mW_0, std=dW_0)
2    pb_0 = compute_interval_probs_bias(np.asarray(vb_0), marg=w_margin, mean=mb_0, std=db_0)
3    pW_1 = compute_interval_probs_weight(np.asarray(vW_1), marg=w_margin, mean=mW_1, std=dW_1)
4    pb_1 = compute_interval_probs_bias(np.asarray(vb_1), marg=w_margin, mean=mb_1, std=db_1)
@@
6    # Now that we have all of the probabilities we just need to multiply them out to get
7    # the final lower bound on the probability of the condition holding.
8    # Work with these probabilities in log space
9    p = 0.0
10    for i in pW_0.flatten():
@@
18    #print math.exp(p)
19    return math.exp(p)

### 28 Feb 2024 v2

@@ Abstract
We introduce two algorithms for computing tight guarantees on the probabilistic robustness of Bayesian Neural Networks (BNNs). Computing robustness guarantees for BNNs is a significantly more challenging task than verifying the robustness of standard Neural Networks (NNs) because it requires searching the parameters’ space for safe weights. Moreover, tight and complete approaches for the verification of standard NNs, such as those based on Mixed-Integer Linear Programming (MILP), cannot be directly used for the verification of BNNs because of the polynomial terms resulting from the consecutive multiplication of variables encoding the weights. Our algorithms efficiently and effectively search the parameters’ space for safe weights by using iterative expansion and the network’s gradient and can be used with any verification algorithm of choice for BNNs. In addition to proving that our algorithms compute tighter bounds than the SoA, we also evaluate our algorithms against the SoA on standard benchmarks, such as MNIST and CIFAR10, showing that our algorithms compute bounds up to 40% tighter than the ${}^{\dagger}$
@@ 1 Introduction
footnotetext: $*{}^{*}$qual contribution.footnotetext: ${}^{\dagger}$he code is available at https://github.com/benbatten/
TightVerificationOfProbabilisticRobustnessInBNNs
@@ Acknowledgements
The authors are grateful to Peyman Hosseini for extensive comments on previous versions of this paper and help with the experiments. This work is supported by the UKRI Centre for Doctoral Training in Safe and Trusted Artificial Intelligence, the “REXASI-PRO” H-EU project, call HORIZON-CL4-2021-HUMAN-01-01, Grant agreement ID: 101070028, and the UK Royal Academy of Engineering (CiET17/18-26).
@@ Appendix A Limitations
1pW_0 = compute_interval_probs_weight(np.asarray(vW_0), marg=w_margin, mean=mW_0, std=dW_0)
2pb_0 = compute_interval_probs_bias(np.asarray(vb_0), marg=w_margin, mean=mb_0, std=db_0)
3pW_1 = compute_interval_probs_weight(np.asarray(vW_1), marg=w_margin, mean=mW_1, std=dW_1)
4pb_1 = compute_interval_probs_bias(np.asarray(vb_1), marg=w_margin, mean=mb_1, std=db_1)
@@
6# Now that we have all of the probabilities we just need to multiply them out to get
7# the final lower bound on the probability of the condition holding.
8# Work with these probabilities in log space
9p = 0.0
10for i in pW_0.flatten():
@@
18#print math.exp(p)
19return math.exp(p)